OpenStudy (anonymous):

How to find the derivative of y=(x^2-3)^2/sqr 4x+5

OpenStudy (anonymous):

you know the quotient law?

OpenStudy (anonymous):

ok, what steps have you taken to try and solve this so far?

OpenStudy (anonymous):

chain and quotient rule. But i'm having problems with the factoring thats all, not the rules

OpenStudy (anonymous):

ok, well let me run through it really quickly

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

the book fully factors the numerator of the answer?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[y=\frac{(x^2-3)^2}{\sqrt{4x+5}}\\y'=\frac{2(2x)(x^2-3)\sqrt{4x+5}-(x^2-3)^2(\frac{1}{2})(4x+5)^{\frac{-1}{2}}(4)}{4x+5}\\=4(4x+5)^{\frac{-1}{2}}\frac{2x(x^2-3)(4x+5)-(x^2-3)}{2(4x+5)}\\=2\frac{8x^4+10x^3-25x^2-30x-3}{(4x+5)^{\frac{3}{2}}}\]

OpenStudy (anonymous):

wow that took 7 minutes. anyways that's as far as id go... really hope I didn't make a mistake lol

OpenStudy (anonymous):

when you multiply by the bottom will it be -1/2 or 1/2 since its in the denominator

OpenStudy (anonymous):

+1/2

OpenStudy (anonymous):

is that the answer in the book?

OpenStudy (anonymous):

the book has 2(x^2-3)(7x^2+10x+3) over the same denominator you have

OpenStudy (anonymous):

hm I made an algebraic error somewhere, but look at the second last line, where I have (x^2-3) as a factor?

OpenStudy (anonymous):

i don't see a mistake. but in the second last line, you factor out the (x^2-3) and you are left with \[y'=\frac{2(x^2-3)(8x^2+10x-1)}{(4x+5)^{\frac{3}{2}}}\]

OpenStudy (anonymous):

ummm can you explain what you did when you factored out the sqr of 4x+5, because i'm still confused with that. Shouldn't the 4x+5 in the equation now have a -1 as an exponent ?

OpenStudy (anonymous):

so when you have \[5a+2a^{-1}\\=a^{-1}(5a^2+2)\] you take out the one with the lowest exponent value. so in this example, its a^(-1) and just think about the exponent laws, and what the exponents inside the brackets shoulw be so that when you multiply the a^(-1) back in hen you get the original EQ again

OpenStudy (anonymous):

but in the first part of the equation you take out a -1/2 and are left with a positive 1 as an exponent, but if you multiply it back you will get a negative 1/2, which isn't what the original was

OpenStudy (anonymous):

ah but then you multiply the (4x+5)^-0.5 by the -(x^2-3) and are back to the first line

OpenStudy (anonymous):

genius, do you see a mistake in my math?

OpenStudy (anonymous):

\[(4x+5)^{\frac{-1}{2}}[(4x+5)-1]\\=(4x+5)^{\frac{1}{2}}-(4x+5)^{\frac{-1}{2}}\]

OpenStudy (anonymous):

0213 you get what i did?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[y=\frac{ (x^2-3)^2 }{ \sqrt{4x+5} }\]\[y'=\frac{ 2(x^2-3)(2x)(\sqrt{4x+5})-\frac{ 1 }{ 2 }(4x+5)^{-\frac{ 1 }{ 2 }}(4)(x^2-3)^2 }{ 4x+5 }\]\[=\frac{ (4x^3-12x)(\sqrt{4x+5})-\frac{ (4)(x^2-3)^2 }{ 2\sqrt{4x+5} } }{ 4x+5 }=\frac{ \frac{ 2\sqrt{4x+5}(4x^3-12x)(\sqrt{4x+5}) }{ 2\sqrt{4x+5} }-\frac{ (4)(x^2-3)^2 }{ 2\sqrt{4x+5} } }{ 4x+5 }\]\[=\frac{ \frac{ 2(4x+5)(4x^3-12x) }{ 2\sqrt{4x+5} }-\frac{ (4)(x^2-3)^2 }{ 2\sqrt{4x+5} } }{ 4x+5 }=\frac{ \frac{ [2(4x+5)(4x^3-12x)]-[4(x^2-3)^2] }{ 2\sqrt{4x+5} } }{ 4x+5 }\]\[=\frac{ \frac{ [32x^4-96x^2+40x^3-120x]-[4(x^4-6x^2+9)] }{ 2\sqrt{4x+5} } }{ 4x+5 }=\frac{ \frac{ [32x^4-96x^2+40x^3-120x]-[4x^4-24x^2+36] }{ 2\sqrt{4x+5} } }{ 4x+5 }\]\[=\frac{ \frac{ 32x^4-96x^2+40x^3-120x-4x^4+24x^2-36 }{ 2\sqrt{4x+5} } }{ 4x+5 }=\frac{ \frac{ 28x^4-72x^2+40x^3-120x-36 }{ 2\sqrt{4x+5} } }{ 4x+5 }\]\[=\frac{ \frac{4( 7x^4-18x^2+10x^3-30x-9) }{ 2\sqrt{4x+5} } }{ 4x+5 }=\frac{ \frac{2( 7x^4-18x^2+10x^3-30x-9) }{ \sqrt{4x+5} } }{ 4x+5 }\]\[= \frac{2( 7x^4-18x^2+10x^3-30x-9) }{ \sqrt{4x+5} } \div \frac{ 4x+5 }{ 1 }\]\[= \frac{2( 7x^4+10x^3-18x^2-30x-9) }{ \sqrt{4x+5} } \times \frac{ 1 }{ 4x+5 }=\frac{2( 7x^4+10x^3-18x^2-30x-9) }{ \sqrt{4x+5}(4x+5) } \]\[=\frac{2( 7x^4+10x^3-18x^2-30x-9) }{ (4x+5)^{\frac{ 3 }{ 2 }} } \rightarrow y'=\frac{2( 7x^4+10x^3-18x^2-30x-9) }{ \sqrt[]{(4x+5)^3} }\] @0213 Took me forever. Hope this helps.

OpenStudy (anonymous):

@matineesuxxx

OpenStudy (anonymous):

omg thank you.

OpenStudy (anonymous):

lol ya I skipped a few steps. thanks :)

OpenStudy (anonymous):

It looks so damn beautiful. I'm gonna keep looking at that the whole day lol.

OpenStudy (anonymous):

haha and that does factor out to what he needed

OpenStudy (anonymous):

What do you mean factor out? The final derivative is in the last row after y'=

OpenStudy (anonymous):

ya but his books factors the numerator

OpenStudy (anonymous):

No, if you factor the top, nothing cancels out. So it's best to leave it unfactored like that. You could factor it too, but that wouldn't make much of a difference.

OpenStudy (anonymous):

btw 0213, I doubt a test would want you to go that in depth or much even ask you to factor. cause like genius said it doesn't make a difference, they are just testing your algebra if they ask you to factor it